how to solve 1=2 by algera

Grade:Select Grade

1 Answers

Saurabh Kumar
8 years ago
Although it is not possible but a kind of joke..

Proof" that 1 + 1 = 1
Let, a = 1
b = 1
a = b
a2 = b2
a2 - b2 = 0
(a-b)(a+b) = 0
(a-b)(a+b)/(a-b) = 0/(a-b)
1(a+b) = 0
(a+b) = 0
1 + 1 = 0
2 = 0
1 = 0
1 + 1 = 1 ….….…....>>>HAPPY